# Improving Phishing Email Detection Performance of Small Large Language Models

### v1

@@ Abstract
Large language models(LLMs) have demonstrated remarkable performance on many natural language processing(NLP) tasks and have been employed in phishing email detection research. However, in current studies, well-performing LLMs typically contain billions or even tens of billions of parameters, requiring enormous computational resources. To reduce computational costs, we investigated the effectiveness of small-parameter LLMs for phishing email detection. These LLMs have around 3 billion parameters and can run on consumer-grade GPUs. However, small LLMs often perform poorly in phishing email detection task. To address these issues, we designed a set of methods including Prompt Engineering, Explanation Augmented Fine-tuning, and Model Ensemble to improve phishing email detection capabilities of small LLMs. We validated the effectiveness of our approach through experiments, significantly improving accuracy on the SpamAssassin dataset from around 0.5 for baseline models like Qwen2.5-1.5B-Instruct to 0.976.
Keywords Phishing Email Detection  $\cdot$ Small LLMs
@@ 1 Introduction
However, we note that although many studies use LLMs for phishing detection, well-performing models typically come with larger parameter counts. For example, Ngoc Tuong Vy Nguyen et al. used a debate framework built with LLaMA-3.1-70b and GPT-4 for phishing detection[11], achieving approximately 0.99 accuracy on datasets like SpamAssassin and Nigerian Fraud[12]. Despite their high accuracy, considering the hundreds of billions or even trillions of parameters in models like LLaMA-3.1-70b[13] and GPT-4, the computational costs are substantial. To address this issue, we propose using smaller language models such as LLaMA-3.2-3B-Instruct[14] and Phi-4-mini-Instruct[15] for detection. These models have around 3 billion parameters, can run on consumer-grade GPUs, and significantly reduce computational requirements. However, we found that directly using these models for detection yields poor results—for instance, Qwen2.5-1.5B-Instruct[16] achieved only 0.388 accuracy on the SpamAssassin dataset, while LLaMA-3.2-3B reached just 0.587. To address these challenges, we propose several methods to enhance the performance of small-parameter LLMs in phishing email detection tasks. By practicing our methods step by step, we effectively improved the phishing email detection ability of small LLMs. The process of our method is shown in Figure 1.
@@
  • Explanation Augmented Fine-tuning: Fine-tuning can enhance LLMs’ capabilities on specific downstream tasks. However, we found that directly fine-tuning on datasets consisting only of emails and their labels yields poor results. We believe that using only email labels as training targets during fine-tuning doesn’t align with the generative nature of LLMs. Therefore, we used PT-4o-mini[17] to generate an explanation for each training sample, detailing why the email is (or is not) a phishing attempt. Through this approach, we effectively improved the detection capabilities of fine-tuned LLMs.
  • Model Ensemble: In the previous step, we fine-tuned multiple LLMs, allowing us to consider model ensemble techniques to further improve detection performance. We implemented two ensemble methods: (1) Majority Vote, where we have three LLMs evaluate each email and take the majority decision as the final result; (2) Confidence Ensemble, where we calculate confidence scores based on the token logprobs from LLM outputs and select the answer with the highest confidence score. Through model ensemble techniques, we further enhanced our detection performance.
@@ 2 Related Work
Despite the outstanding performance of LLMs in phishing email detection tasks, a critical challenge remains: the large number of parameters imposes significant obstacles to model deployment and costly API call expenses. Additionally, fine-tuning these large LLMs for downstream tasks demands substantial computational and time resources. Our study focuses on evaluating the performance of small LLMs in phishing email detection. Through fine-tuning, these LLMs can achieve performance that closely approaches that of large LLMs, which makes them an excellent choice for balancing cost and effectiveness while preserving interpretability.
@@ 3.3 Explanation Augmented Fine-tuning
$h=Wx+\Delta Wx=Wx+\frac{\alpha}{r}ABx$
@@ 3.4 Model Ensemble
In the forward pass, the final logits pass through a softmax layer to obtain probabilities corresponding to different tokens, which we can obtain in the form of logprobs. The sequence probability is calculated by exponentiating the logprobs of each token multiplying these probabilities together, and then taking the $N$th root of the result, where $N$ represents the sequence length. This process effectively computes the geometric mean of token probabilities, providing a Length-Normalized(LN) confidence score.
@@
$LN~{}Conf.=\left(\prod_{i=1}^{N}\exp(logprob_{i})\right)^{1/N}$
4 Experiment
In this section, we conduct experiments to demonstrate the effectiveness of our method. We use the SpanAssassin dataset for our experiment, and the evaluation was conducted on a held-out test set consisting of 1,069 emails. The small LLMs employed include LLaMA-3.2-3B-Instruct, Phi-4-mini-Instruct, and Qwen-2.5-1.5B-Instruct, with parameters of 3 billion, 3.8 billion, and 1.5 billion, respectively. Both inference and fine-tuning can be performed on one RTX 3090 GPU, and we limit the length of training sequences and output sequences to avoid GPU memory overflow issues. We first conduct zero-shot experiments based solely on prompt engineering, followed by explanation augmented fine-tuning and model ensemble experiments. We also perform ablation studies to illustrate the significant performance improvements achieved through explanation enhancement in the training data.
Evaluation Metrics: We report standard classification metrics including Precision, Recall, and F1-Score, computed as:
@@ 4.1 Vanilla Prompt Engineering Performance
Model Accuracy F1 Precision Recall
LLaMA-3.2-3B-Instruct 0.587 0.543 0.934 0.922
Phi-4-mini 0.647 0.537 0.925 0.963
Qwen-2.5-1.5b-Instruct 0.388 0.469 0.863 0.551
From the result, we can observe that: (1) Applying Vanilla Prompting on small models performs very poorly, with accuracy lower than 0.7, and F1 scores around 0.5; (2) Models with smaller parameter counts perform worse, with the Qwen 1.5B model achieving only 0.388 accuracy, which is 40% lower than Phi-4-mini (3.8B).
@@ 4.2 Explanation Augmented Fine-tuning Performance
We then perform augmented fine-tuning on selected LLMs. We extract a subset of 1,000 samples from the SpamAssassin training set and perform explanation augmentation on this subset. We then fine-tune these LLMs with this training set using LoRA. We compare the results of our fine-tuned small LLMs with traditional ML models such as Naive Bayes, Support Vector Machine (SVM), and XGBoost, as along with GPT-3.5-Turbo, GPT-4o-mini and LLaMA-3.1-70B-Instruct. When using traditional ML, we employed the paraphrase-MiniLM-L3-v2 embedding model from Sentence-BERT to embed emails, and used the resulting embedding vectors for training and testing. The results are shown in Table 2.
After explanation augmented fine-tuning, the performance of small LLMs on the SpamAssassin dataset improved significantly. The accuracy and F1 score of LLaMA-3.2-3B-Instruct model increased from 0.587 and 0.543 to 0.963 and 0.928, respectively, while Phi-4-mini’s increased from 0.647 and 0.537 to 0.968 and 0.944. Qwen-2.5-1.5b-Instruct’s accuracy improved by 122%, from 0.388 to 0.860. Additionally, these small LLMs even outperformed larger LLMs after fine-tuning. For example, Phi-4-mini’s accuracy was 27.8% higher than GPT-3.5-Turbo, 21.8% higher than LLaMA-3.1-70B-Instruct, and 2.1% higher than GPT-4o-mini. The fine-tuned LLaMA-3.2-3B-Instruct also achieved better accuracy and F1 scores than these three large LLMs. The performance of small LLMs was similar to XGBoost and SVM, but they can provide interpretable judgment results in the form of human language. Our experimental results demonstrate that post-fine-tuning, small LLMs exhibit substantially enhanced phishing email detection capabilities, achieving performance metrics comparable to (in some cases exceeding) those of established machine learning approaches. Notably, these fine-tuned compact models outperform larger LLMs with orders of magnitude more parameters, highlighting the effectiveness of our approach in optimizing model efficiency for detecting phishing emails.
@@
Table 2: The experimental results of GPT-3.5-Turbo, GPT-4o-mini, traditional ML models and Explanation Augmented Fine-tuning. EA represents ’Explanation Augmented Fine-tuned’
Model Accuracy F1 Precision Recall
Naive Bayes 0.852 0.760 0.740 0.780
SVM 0.967 0.945 0.950 0.940
XGBoost 0.953 0.920 0.941 0.900
GPT-3.5-Turbo 0.757 0.705 0.553 0.972
GPT-4o-mini 0.948 0.915 0.941 0.891
LLaMA-3.1-70B-Instruct 0.801 0.635 0.693 0.586
LLaMA-3.2-3B-Instruct-EA 0.963 0.928 0.934 0.922
Phi-4-mini-EA 0.968 0.944 0.925 0.963
Qwen-2.5-1.5b-Instruct-EA 0.860 0.673 0.863 0.551
@@ 4.3 Model Ensemble
After fine-tuning multiple LLMs, we ensemble these models to fully leverage the knowledge and capabilities of different LLMs. We employed confidence ensemble and majority vote methods, and compared the results of these ensemble approaches with those of only fine-tuned models, as shown in Table 3. Note that for confidence ensemble, we only use LLaMA-3.2-3B-Instruct and Phi-4-mini, while for majority vote, we include all three small LLMs.
The experimental results show that confidence ensemble and majority vote further improved the accuracy to around 0.975, with F1 scores reaching 0.953 and 0.959, respectively. Through model ensemble, we further improved the capability and stability of LLMs in phishing email detection. However, the improvement margin was relatively limited compared to the gains achieved through fine-tuning.
Table 3: The experimental results of Confidence Ensemble and Majority Vote.
Model Accuracy F1 Precision Recall
LLaMA-3.2-3B-Instruct-EA 0.963 0.928 0.934 0.922
Phi-4-mini-EA 0.968 0.944 0.925 0.963
Qwen-2.5-1.5b-Instruct-EA 0.860 0.673 0.863 0.551
Confidence Ensemble 0.975 0.953 0.958 0.947
Majority Vote 0.976 0.959 0.980 0.937
@@ 4.4 Ablation Study
Our experiments reveal that without explanations in the fine-tuning data, the final prediction results are significantly compromised. The accuracy of the LLaMA-3.2-3B-Instruct model decreased by 40.7%, while its F1 score plummeted from 0.928 to 0.219. The accuracy and F1 score of Qwen-2.5-1.5B-Instruct decreased by 36.4% and 26.0%, respectively. Phi-4-mini demonstrated relatively robust performance, but its accuracy and F1 score still decreased by 15.1% and 27.1%, respectively. Overall, through ablation studies, we found that explanation augmentation has a significant impact on fine-tuning effectiveness, as it transforms our phishing email detection dataset from a closed-form prediction dataset to an open-ended generative dataset.
@@
Table 4: The experimental results of using and not using Explanation Augmented Fine-tuning. EA represents ’Explanation Augmented Fine-tuned’
Model Accuracy F1 Precision Recall
LLaMA-3.2-3B-Instruct-EA 0.963 0.928 0.934 0.922
Phi-4-mini-EA 0.968 0.944 0.925 0.963
Qwen-2.5-1.5b-Instruct-EA 0.860 0.673 0.863 0.551
LLaMA-3.2-3B-Instruct-no-EA 0.571 0.219 0.221 0.216
Phi-4-mini-EA-no-EA 0.822 0.688 0.621 0.771
Qwen-2.5-1.5b-Instruct-no-EA 0.547 0.498 0.376 0.735

### v2

@@ Abstract
Large language models(LLMs) have demonstrated remarkable performance on many natural language processing(NLP) tasks and have been employed in phishing email detection research. However, in current studies, well-performing LLMs typically contain billions or even tens of billions of parameters, requiring enormous computational resources. To reduce computational costs, we investigated the effectiveness of small-parameter LLMs for phishing email detection. These LLMs have around 3 billion parameters and can run on consumer-grade GPUs. However, small LLMs often perform poorly in phishing email detection task. To address these issues, we designed a set of methods including Prompt Engineering, Explanation Augmented Fine-tuning, and Model Ensemble to improve phishing email detection capabilities of small LLMs. We validated the effectiveness of our approach through experiments, significantly improving both accuracy and F1 score on the SpamAssassin and CEAS_08 datasets. Furthermore, the fine-tuned models demonstrated strong transferability, achieving robust performance across multiple unseen phishing datasets, outperforming traditional baselines and approaching standard-sized LLMs.
Keywords Phishing Email Detection  $\cdot$ LLMs  $\cdot$ Fine-tuning
@@ 1 Introduction
However, we note that although many studies proposed using LLMs for phishing detection, well-performing models typically come with larger parameter counts. For example, Ngoc Tuong Vy Nguyen et al. used a debate framework built with LLaMA-3.1-70B and GPT-4 for phishing detection[11], achieving approximately 0.99 accuracy on datasets like SpamAssassin and Nigerian Fraud[12]. Despite their high accuracy, considering the hundreds of billions or even trillions of parameters in models like LLaMA-3.1-70B[13] and GPT-4[10], the computational costs are substantial. To address this issue, we propose using smaller language models such as LLaMA-3.2-3B-Instruct[14] and Phi-4-mini-Instruct[15] for detection. These models have around 3 billion parameters, can run on consumer-grade GPUs, and significantly reduce computational requirements. Furthermore, these smaller LLMs can provide text-level evidence for their judgments, thereby offering a degree of interpretability to the results. However, we found that directly using these models for detection yields poor results (e.g. strong bias or inconsistent output formats). To address these challenges, we propose several methods to enhance the performance of small-parameter LLMs in phishing email detection tasks. By practicing our methods step by step, we effectively improved the phishing email detection ability of small LLMs. The process of our method is shown in Figure 1.
@@
  • Explanation Augmented Fine-tuning: Fine-tuning can enhance LLMs’ capabilities on specific downstream tasks. However, we found that directly fine-tuning on datasets consisting only of emails and their labels yields poor results. We believe that using only email labels as training targets during fine-tuning doesn’t align with the generative nature of LLMs. Therefore, we used GPT-4o-mini[16] to generate an explanation for each training sample, detailing why the email is (or is not) a phishing attempt. Then, we fine-tuned each LLM on datasets consisting of emails, labels and explanations. Through this approach, we effectively improved the detection capabilities of fine-tuned LLMs.
  • Model Ensemble: In the previous step, we fine-tuned multiple LLMs, allowing us to consider model ensemble techniques to further improve detection performance. We implemented two ensemble methods: (1) Majority Vote, where we have three LLMs evaluate each email and take the majority decision as the final result; (2) Confidence Ensemble, where we calculate confidence scores based on the token logprobs from LLM outputs and select the answer with the highest confidence score. Through model ensemble techniques, we found that it is possible to further enhance detection performance.
@@ 2 Related Work
Despite the outstanding performance of LLMs in phishing email detection tasks, a critical challenge remains: the large number of parameters imposes significant obstacles to model deployment and costly API call expenses. Additionally, fine-tuning these large LLMs for downstream tasks demands substantial computational and time resources. Our study focuses on evaluating the performance of small LLMs in phishing email detection. Through fine-tuning, these smaller LLMs can achieve performance that meets or even surpasses that of large LLMs, making them an excellent choice for balancing cost and effectiveness while preserving interpretability.
@@ 3.3 Explanation Augmented Fine-tuning
$h=Wx+\Delta Wx=Wx+\frac{\alpha}{r}ABx$ (2)
@@ 3.4 Model Ensemble
In the forward pass, the final logits pass through a softmax layer to obtain probabilities corresponding to different tokens, which we can obtain in the form of logprobs. The sequence probability is calculated by exponentiating the logprobs of each token multiplying these probabilities together, and then taking the $N$th root of the result, where $N$ represents the sequence length. This process effectively computes the geometric mean of token probabilities, providing a Length-Normalized(LN) confidence score, see Equation 3.
@@
$LN~{}Conf.=\left(\prod_{i=1}^{N}\exp(logprob_{i})\right)^{1/N}$ (3)
4 Experiments
In this section, we conduct experiments to demonstrate the effectiveness of our method. The small LLMs employed include LLaMA-3.2-3B-Instruct[14], Phi-4-mini-Instruct[15], and Qwen-2.5-1.5B-Instruct[28], with parameters of 3 billion, 3.8 billion, and 1.5 billion, respectively. We fine-tuned the three LLMs separately on SpamAssassin and CEAS_08. Each model was fine-tuned and evaluated separately on each of the two datasets. Specifically, for each dataset, a model was fine-tuned on its training split and subsequently evaluated on its corresponding test split. Both inference and fine-tuning can be performed on one RTX 3090 GPU, and we limit the length of training sequences and output sequences to avoid GPU memory overflow issues. We first conduct zero-shot experiments based solely on prompt engineering, followed by explanation-augmented fine-tuning and model ensemble experiments. We also perform ablation studies to illustrate the significant performance improvements achieved through explanation enhancement in the training data. Moreover, we evaluate the transferability of our fine-tuned LLMs compared with traditional models and standard-sized LLMs. There are 1,069 test samples in the SpamAssassin test set, and 1,000 test samples each in the CEAS_08, Ling, and Enron test sets.
Evaluation Metrics: We report standard classification metrics including Precision, Recall, and F1 score, computed as:
@@ 4.1 Vanilla Prompt Engineering Performance
Model SpamAssassin CEAS_08
Accuracy F1 Precision Recall Accuracy F1 Precision Recall
LLaMA-3.2-3B-Instruct 0.587 0.354 0.317 0.401 0.671 0.566 0.591 0.543
Phi-4-mini-Instruct 0.647 0.537 0.423 0.734 0.808 0.798 0.804 0.793
Qwen-2.5-1.5B-Instruct 0.388 0.469 0.899 0.317 0.657 0.758 0.641 0.927
From the results, we observe that applying Vanilla Prompting to small LLMs yields unsatisfactory performance. On the SpamAssassin dataset, accuracies are below 0.65 and F1-scores hover around 0.5. Although performance on CEAS_08 is relatively better, the best accuracy and F1-score are only around 0.8. These results demonstrate that small LLMs without fine-tuning are inadequate for phishing email detection tasks (especially in high-risk environments), underscoring the need for enhancement methods such as fine-tuning.
@@ 4.2 Explanation Augmented Fine-tuning Performance
Table 2: The experimental results of GPT-3.5-Turbo, GPT-4o-mini, traditional ML models and EA(Explanation Augmented Fine-tuned) LLMs
Model SpamAssassin CEAS_08
Accuracy F1 Precision Recall Accuracy F1 Precision Recall
Naive Bayes 0.852 0.760 0.740 0.780 0.829 0.835 0.947 0.746
SVM 0.967 0.945 0.950 0.940 0.989 0.991 0.990 0.991
XGBoost 0.953 0.920 0.941 0.900 0.997 0.997 0.998 0.997
GPT-3.5-Turbo 0.757 0.705 0.553 0.972 0.880 0.897 0.882 0.912
GPT-4o-mini 0.948 0.915 0.941 0.891 0.854 0.859 0.966 0.773
LLaMA-3.1-70B-Instruct 0.801 0.635 0.693 0.586 0.881 0.896 0.882 0.910
LLaMA-3.2-3B-Instruct-EA 0.963 0.928 0.934 0.922 0.891 0.893 0.955 0.838
Phi-4-mini-Instruct-EA 0.968 0.944 0.925 0.963 0.917 0.927 0.932 0.922
Qwen-2.5-1.5B-Instruct-EA 0.860 0.673 0.863 0.551 0.860 0.868 0.907 0.833
We then perform augmented fine-tuning on selected LLMs using the SpamAssassin and CEAS_08 datasets respectively. We extract a subset of 1,000 samples from the training set and perform explanation augmentation on this subset. We then fine-tune these LLMs with this training set using LoRA. We compare the results of our fine-tuned small LLMs with traditional ML models such as Naive Bayes, Support Vector Machine (SVM), and XGBoost, as along with standard-sized LLMs such as GPT-3.5-Turbo, GPT-4o-mini and LLaMA-3.1-70B-Instruct. When using traditional ML, we employed the paraphrase-MiniLM-L3-v2 embedding model from Sentence-BERT to embed email subject and body, and used the resulting embedding vectors for training and testing[29]. Unlike LLMs, traditional models are trained without explanation augmentation, because they are not designed for open-ended generative tasks. Thus, their training sets only include email content and labels, without any explanation. The results are shown in Table 2 .
After explanation augmented fine-tuning, the performance of small LLMs improved significantly. On the SpamAssassin dataset, the accuracy and F1 score of LLaMA-3.2-3B-Instruct model increased from 0.587 and 0.543 to 0.963 and 0.928, respectively, while Phi-4-mini-Instruct’s increased from 0.647 and 0.537 to 0.968 and 0.944. Qwen-2.5-1.5B-Instruct’s accuracy improved by 122%, from 0.388 to 0.860. On the CEAS_08 dataset, the LLaMA-3.2-3B-Instruct model achieved an accuracy of 0.891 and an F1 score of 0.893, while Phi-4-mini-Instruct reached 0.917 and 0.927, respectively. Although these results are marginally lower than their corresponding performance on the SpamAssassin dataset, they nonetheless demonstrate a substantial improvement over their vanilla baselines. Remarkably, Qwen-2.5-1.5B-Instruct attained an accuracy of 0.860 and an F1 score of 0.868, outperforming its performance on the SpamAssassin dataset. Although the performance of Qwen-2.5-1.5B was not as good as LLaMA-3.2-3B and Phi-4-mini-Instruct, due to its smaller parameter counts, it still showed substantial improvement over its own vanilla baseline. Additionally, these small LLMs even outperformed larger LLMs after fine-tuning. For example, on the SpamAssassin dataset, Phi-4-mini-Instruct’s accuracy was 27.9% higher than GPT-3.5-Turbo, 20.8% higher than LLaMA-3.1-70B-Instruct, and 2.1% higher than GPT-4o-mini. The fine-tuned LLaMA-3.2-3B-Instruct also achieved better accuracy and F1 scores than these three large LLMs. On the CEAS_08 dataset, both Phi-4-mini-Instruct and LLaMA-3.2-3B-Instruct surpassed larger-scale LLMs. Among them, Phi-4-mini-Instruct delivered the most competitive results, achieving an accuracy 4.2% higher than GPT-3.5-Turbo, 7.4% higher than GPT-4o-mini, and 4.1% higher than LLaMA-3.1-70B-Instruct. The performance of small LLMs was similar to XGBoost and SVM, but they can provide interpretable judgment results in the form of human language.
Our experimental results demonstrate that after fine-tuning, small LLMs exhibit substantially enhanced phishing email detection capabilities, achieving performance metrics comparable to (in some cases exceeding) those of established machine learning approaches. Notably, these fine-tuned compact models outperform larger LLMs with orders of magnitude more parameters, highlighting the effectiveness of our approach in optimizing model efficiency for detecting phishing emails.
@@ 4.3 Model Ensemble
After fine-tuning multiple LLMs, we explored model ensemble strategies to potentially leverage the complementary strengths of different LLMs. We implemented two ensemble methods: confidence ensemble and majority vote. The comparative results between ensemble approaches and individual fine-tuned models are presented in Table 3. For the confidence ensemble, we specifically utilized LLaMA-3.2-3B-Instruct and Phi-4-mini-Instruct, while the majority vote incorporated all three fine-tuned small LLMs.
The experimental results demonstrate that on the SpamAssassin dataset, both ensemble methods achieved notable improvements, with accuracies reaching approximately 0.975 and F1 scores of 0.953 and 0.959 for confidence ensemble and majority vote, respectively. However, on the CEAS_08 dataset, neither ensemble method significantly outperformed the best individual model (Phi-4-mini-Instruct), suggesting that the ensemble benefits may vary across different datasets.
While model ensemble strategies demonstrated potential in enhancing the robustness and performance of LLMs in phishing email detection, the marginal improvements were relatively modest compared to the substantial gains achieved through explanation-augmented fine-tuning alone.
Table 3: The experimental results of Confidence Ensemble and Majority Vote comparing to using EA(Explanation Augmented Fine-tuned) small LLMs
Model SpamAssassin CEAS_08
Accuracy F1 Accuracy F1
LLaMA-3.2-3B-Instruct-EA 0.963 0.928 0.891 0.893
Phi-4-mini-Instruct-EA 0.968 0.944 0.917 0.927
Qwen-2.5-1.5B-Instruct-EA 0.860 0.673 0.860 0.868
Confidence Ensemble 0.975 0.953 0.895 0.906
Majority Vote 0.976 0.959 0.884 0.896
@@ 4.4 Ablation Study
Our experiments reveal that without explanations in the fine-tuning data, the final prediction results are significantly compromised. The accuracy of the LLaMA-3.2-3B-Instruct model decreased by 40.7%, while its F1 score plummeted from 0.928 to 0.219. The accuracy and F1 score of Qwen-2.5-1.5B-Instruct decreased by 36.4% and 26.0%, respectively. Phi-4-mini-Instruct demonstrated relatively robust performance, but its accuracy and F1 score still decreased by 15.1% and 27.1%, respectively. Overall, through ablation studies, we found that explanation augmentation has a significant impact on fine-tuning effectiveness, as it transforms our phishing email detection dataset from a closed-form prediction dataset to an open-ended generative dataset.
@@
Table 4: The experimental results of using and not using EA(Explanation Augmented Fine-tuned) small LLMs
Model SpamAssassin CEAS_08
Accuracy F1 Precision Recall Accuracy F1 Precision Recall
LLaMA-3.2-3B-Instruct-EA 0.963 0.928 0.934 0.922 0.891 0.893 0.955 0.838
Phi-4-mini-Instruct-EA 0.968 0.944 0.925 0.963 0.917 0.927 0.932 0.922
Qwen-2.5-1.5B-Instruct-EA 0.860 0.673 0.863 0.551 0.860 0.868 0.907 0.833
LLaMA-3.2-3B-Instruct-no-EA 0.571 0.219 0.221 0.216 0.346 0.147 0.158 0.137
Phi-4-mini-Instruct-no-EA 0.822 0.688 0.621 0.771 0.814 0.799 0.760 0.842
Qwen-2.5-1.5B-Instruct-no-EA 0.547 0.498 0.376 0.735 0.573 0.540 0.545 0.534
4.5 Experiment on Transferability
This section evaluates the transferability of our fine-tuned models alongside other methods. To simulate real-world generalization challenges, we trained small LLMs and traditional machine learning models (such as SVM, XGBoost and Naive Bayes) on a source dataset (either CEAS_08 or SpamAssassin) and subsequently tested their performance on unseen datasets: Enron and Ling[12]. For broader comparison, we also include the results of large LLMs like GPT-3.5-Turbo. We report the accuracy and F1 score to compare how well each model generalizes across domains as shown in Fig. 3 and Fig. 4.
Trained on SpamAssassin: After fine-tuning on the SpamAssassin dataset, both small LLMs and traditional models (such as SVM) demonstrated excellent performance on this dataset. Except for the Qwen-2.5-1.5B-Instruct model, all models achieved approximately 0.95 in both accuracy and F1 scores, with Phi-4-mini-Instruct and LLaMA-3.2-3B-Instruct slightly outperforming traditional models like SVM. Small LLMs and traditional models also exhibited outstanding performance on the Ling dataset. Despite not being trained on the corresponding datasets, the explanation-enhanced fine-tuned Phi-4-mini-Instruct and LLaMA-3.2-3B-Instruct models achieved accuracies of 0.972 and 0.978, and F1 scores of 0.906 and 0.923, respectively. SVM and XGBoost models also attained high accuracies of 0.938 and 0.924, with F1 scores reaching 0.815 and 0.753.
However, on the CEAS_08 and Enron datasets, small LLMs demonstrated notably better transferability compared to traditional models. Phi-4-mini-Instruct maintained an accuracy of 0.858 on the CEAS_08 dataset, surpassing SVM and XGBoost by 4.8% and 14.1% respectively, while also achieving higher F1 scores. On the Enron dataset, Phi-4-mini-Instruct, LLaMA-3.2-3B-Instruct, and Qwen-2.5-1.5B-Instruct achieved accuracies of 0.908, 0.878, and 0.771 respectively, outperforming Naive Bayes, SVM, and XGBoost which scored 0.597, 0.716, and 0.644. Small LLMs generally achieved higher F1 scores as well.
Additionally, it is noteworthy that Phi-4-mini-Instruct and LLaMA-3.2-3B-Instruct outperformed the three regular-sized LLMs on all datasets except CEAS_08, although the number of parameters of the latter is several times or even dozens of times that of the former.
Trained on CEAS_08: After fine-tuning on the CEAS_08 dataset, traditional models such as SVM and XGBoost demonstrated strong performance on the test set, achieving accuracy and F1 scores close to 1. However, their performance dropped dramatically on the other three datasets, with accuracies of only 0.785 and 0.836 on the Ling dataset, and F1 scores below 0.5. The highest accuracy these two models achieved on the remaining datasets was merely 0.804. In contrast, Phi-4-mini-Instruct and LLaMA-3.2-3B-Instruct models maintained accuracies above 0.85 and F1 scores above 0.8 across all four datasets, showing no significant dataset bias. Furthermore, regular-sized LLMs generally performed slightly worse than the fine-tuned small LLMs in most cases.
In summary, small LLMs enhanced through explanation augmented fine-tuning demonstrate superior performance not only on the training dataset but also across various unseen datasets. This empirical evidence strongly indicates that our approach exhibits better transferability compared to traditional machine learning models, while consistently outperforming regular-sized LLMs in detecting phishing emails.
Figure 3: Accuracy and F1 scores of different models evaluated on multiple datasets. Note that small-scale LLMs and traditional machine learning models are specifically trained on the SpamAssassin dataset, while for standard-sized LLMs (e.g., GPT-3.5-Turbo), we employ direct prompting for phishing email detection without additional training.
Figure 4: Accuracy and F1 scores of different models evaluated on multiple datasets. Note that small-scale LLMs and traditional machine learning models are specifically trained on the CEAS_08 dataset, while for standard-sized LLMs (e.g., GPT-3.5-Turbo), we employ direct prompting for phishing email detection without additional training.